What mass of ZnCl2 can be prepared from the reaction of 3.27 grams of zinc with 3.30 grams of HCl

1
Expert's answer
2022-04-07T04:06:49-0400

Moles of zinc = "\dfrac{mass}{RAM} = \dfrac{3.27g}{65.38g/mol}" = 0.05mol


The balanced equation for the reaction


Zn + 2HCl "\to" ZnCl2 + H2


Mole ratios of Zn: HCl is 1:2


Moles of HCl required is therefore;

= "\dfrac{2}{1}x0.05 = 0.1mol"


But there are only = "\dfrac{3.30g}{36.458g/mol}=0.0905 moles of HCl"


Thus HCl is the limiting reagent


Moles of the product is therefore = "\dfrac{1}{2}x0.0905 = 0.045mol of ZnCl2"


Mass or product = moles x MM = 0.045mol x 136.286g/mol

= 6.168g